# The role of turbulence in setting the phase of the ISM and implications for the star formation rate

### v2

@@ 2.6 Simulation suite with constant UV field
The parameters of the first set of simulations are summarised in Tables 1 and the final snapshot of each simulation is displayed in Figure 3. We can see a wide variety in terms of structure. The low-$\Sigma$ runs (top panels), especially the ones with not too weak driving, have a prominent diffuse gas component with a thin disk of denser material in which stars form. The high-$\Sigma$ runs (bottom panels) are more filamentary. Increasing the driving strength leads to more large-scale structure, seen as large overdensities and voids (Colman et al., 2022).
@@ 3.1 Star formation history
The qualitative findings described here for modest column densities are different from the results obtained for columns densities above 20 $\mathrm{M}_{\odot}$$\mathrm{pc}^{-2}$, where large-scale driving was found to always reduces the SFR (Brucy et al., 2020, 2023). This indicates that the relative importance of the processes involved in star formation varies with the column density of the region.
@@ 4.1 Gas phases
In what follows, we will adopt the following definitions for the gas phases. Neutral gas with a temperature below 250 K is considered to be CNM. This threshold temperature is in agreement with the one used in observations (McClure-Griffiths et al., 2023). Note that this definition also includes molecular hydrogen, which is expected to form from cold gas at high densities but whose formation is not tracked explicitly in these simulations. For the lower bound on the temperature of the WNM we choose 3000 K, a value slightly below what is proposed in McClure-Griffiths et al. (2023), to include our lower temperature WNM branch. The LNM is then everything with temperatures in between these thresholds. Our boundaries are shown as dotted lines in Figure 5.
@@ 4.3 Dense gas
We further quantify the efficiency of the formation of dense gas in cold clouds in Figure 8, where we plot the ratio between the amount of CNM versus the amount of gas with densities greater than 100 cm-3. For a point in the bottom left part of the diagram, the amount of cold material is small and the conversion to high density gas inefficient. If a point is in he top part of the diagram, it means the conversion to dense gas is efficient. In the low-$\Sigma$ regime the conversion from cold to dense gas occurs at a roughly equal, albeit low, efficiency which is independent of the level of externally driven turbulence. While the amount of cold HI is similar for all high-$\Sigma$ runs, there is about a factor 2 to 3 less dense gas for the strong driving case. This brings the efficiency to the levels of the lower column density simulations. Strong driving hinders the creation of dense self-gravitating clumps. This is in line with predictions by analytical theories of gravo-turbulent fragmentation (Hennebelle et al., 2024).
@@ 4.4 A dual role of turbulence on the thermal state?
Here we demonstrated that turbulence possibly has a second, thermal, dual nature which typically occurs at larger scales. Hennebelle & Pérault (1999) already showed that turbulence can create cold condensations in the ISM. Here, we saw that it acts as a perturbing force, promoting phase change and moving the gas away from thermal equilibrium, seen by a consistent increase in LNM fraction (Figure 6). The duality then lays in the creation versus destruction of CNM. Our simulation results seem to indicate turbulence has a stronger impact on the WNM-to-CNM conversion than on the reverse conversion. In fact, there is no indication that turbulence is able to bring the gas back into the WNM state, as the WNM fraction decreases and more gas is in the unstable regime. However, we have to keep in mind that turbulence is not only driven by the external large-scale force that is applied during the simulation. Stellar feedback also drives turbulence. For the set of high-$\Sigma$ simulations, the turbulent energy injection by SN is either stronger or comparable to that of the large-scale driving, except for the run with strong driving (Figure 1). Furthermore, the SFR in those three runs is similar, so only in the case of strong driving would we be able to directly link a change in the gas-phase fractions to an excess of turbulence. If the level of turbulence is a determining factor for CNM formation, we would expect the three other simulations to have similar CNM fractions, as is indeed the case. Possibly the stellar feedback self-regulated the WNM and CNM fractions. Controlled simulation of a multi-phase ISM without stellar feedback, could help quantify the dual role of turbulence on the CNM fraction further. Note that the role of feedback is more complex than just injecting turbulence, as massive stars emit UV radiation which contributes directly to the heating of the ISM.
@@ 4.5 Caveats of a uniform UV background
The transition point between WNM and CNM depends on the UV radiation field, metallicity and hydrogen ionisation rate (Wolfire et al., 2003). While our simulations do track the local UV photon field which is used to determine the ionisation of the gas, this information is not used in the cooling model where a uniform background UV field fixed to the solar neighbourhood value $G_{0}$ is assumed. Observations in the Milky-Way and nearby galaxies show that the UV field strength decreases as a function of galactic radius. The UV field in our low-$\Sigma$ runs may thus be too high compared to real low density regions in galaxies. A lower UV field implies less heating which would lead to naturally higher CNM fraction. Furthermore, in the majority of environments we expect the UV field to be directly coupled to the formation of new stars, especially massive ones. In the next section, we present the result from simulations where the uniform UV background is proportional to the SFR according to the recipe of Ostriker et al. (2010). We will see that this indeed result in more structure formation and enhanced SFRs compared to the low-$\Sigma$ runs presented in this section. It remains however unclear how local variations in the UV field may affect the overall CNM fraction. While we demonstrated a clear influence of turbulence under a fixed UV field, in practice regions with conditions similar to those in the low-$\Sigma$ runs may be rare.
@@ 5 Variable UV background
5.1 Simulation suite with variables UV background
@@
A key difference is the UV background, which is variable in time and proportional to the SFR measured in the computational domain. Indeed, we expect the UV field to be proportional to the SFR (Ostriker et al., 2010). The mean FUV density relative to the solar neighbourhood value $G^{\prime}_{0}$ is
@@
Additionally, we set a minimum UV field of 0.1 $G_{0}$. We use the implementation of Brucy et al. (2023).
@@ 5.4 Velocity dispersion
Figure 11 displays the velocity dispersion as a function of time in the eight simulations of Table 2. A comparison with the runs presented in Figure 2 is particularly enlightening. Except for the two cases with the strongest driving, the velocity dispersions in this set of simulations reach about 11-13 km s-1, irrespectively of the forcing strength and initial column density. In contrast, when the UV background is kept constant, the runs with 10.2 M pc-2 present significantly lower velocity dispersion ranging from 2 km s-1 when no turbulent driving is applied, to 7 km s-1 when turbulent forcing operates. This stems from the much higher star formation rate that is achieved in the simulation with variable UV background. The stellar feedback, in particular the number of supernova explosions, is one to two orders of magnitude larger than in the case of constant UV background.
@@ 5.5 Constant or variable UV background?
Although a UV background which is local and proportional to the SFR may be assumed to be the norm, there are situations where a constant UV field may be a good approximation. Let us remind that the mean free path of the UV photons in an ISM with a mean density of 0.3 cm-3 is about 1 kpc. This implies that for instance in the outskirts of galaxies, the ISM is heated by UV photons which have been emitted at several kpc distance. Accumulated over the galaxy and in a time-span considered by our simulations, the background can then be considered fairly constant.
@@
Figure 12: Relations of the SFR with column density for different gas reservoirs. Top: simulation set with constant UV (Table 1), bottom: set with variables UV (Table 2). The SFR is averaged over a time scale of 10 Myr after (full circles) or before (empty circles) the time of the snapshot at which the gas column densities have been measured. In the left panels, the contours sketch the location of the data point from Bigiel et al. (2008): the outer contour tracers roughly all points, while inner one corresponds roughly to their orange contours denoting 2 independent data point per 0.05 dex. In the middle panels, we show a power-law with index 1.4 to guide the eye. In the right panel, we show the linear relation found by Bigiel et al. (2008) between the SFR and the column density of H2 gas, and extrapolate to lower densities. The simulation data is colour-coded as in Figures 6 and 10.
@@ 6.1 SFR versus different gas phases
In the left panels of Figure 12, we show the total neutral gas versus the SFR. We compare with the observational results from Bigiel et al. (2008), who studied the KS-relation in a sample of 18 galaxies. Their results for HI + H2 are shown as contours, where the outer contour sketches the full extend of their data points and the inner contour indicate where most of the measurements are situated. The observations are characterised by a break at 9 $\mathrm{M}_{\odot}$pc-2. For higher column densities, a power-law relation with index 1.4 is found, while for lower column densities the relation is steeper with a lot of scatter (Kennicutt & Evans, 2012). The results from the high-$\Sigma$ runs with constant UV field (top panels) fall in the observed range, though the SFR might be a bit high compared to the majority of the observations. Possibly this is because we do not include early feedback in the form of stellar winds and jets, which are expected to reduce the SFR by a factor 2 (Rathjen et al., 2021; Verliat et al., 2022). At low column density, we reproduce the large scatter. The simulations without driving produce SFRs which are below the observed range. This could mean that the environment modelled by this simulation is unlikely to occur in real galaxies. Indeed, large-scale dynamic effects such as rotation or infall would provide a minimal level of turbulence driving. The simulations with strongest driving (orange and green points) generally agree best with the observations.
@@
The middle panels shows the relation with cold gas, that is CNM and cold dense gas which would be molecular. A positive correlation now appears at low column densities. These findings are in line with the results from Smith et al. (2023) who studied the distribution of CNM in their full galaxy simulation. Overall, the relation between the column density of cold gas and the SFR exhibits a scatter of about one order of magnitude around a power-law with index 1.4. This is the same index as found for the relation with total gas, above a column density of 20 $\mathrm{M}_{\odot}$ pc-2. This leads us to the conclusion that the scatter in the observations of the KS-relation at low column densities, as shown in the left panel, is likely due to regional differences in the CNM fraction. Following this line of thought, it implies that at higher column densities the CNM fraction on a scale of 1 kpc does not vary significantly between regions. Indeed, in the simulations we find it to be around 40 % independent of the external driving strength.
Also for the simulations with variables UV background in the bottom panel, we find a similar correlation between the SFR and the cold gas. Again, the strong driving run deviates in behaviour: there is a large variation in SFR for similar column densities. The large velocity dispersions in these runs (see Figure 11), may prevent the formation of sink-forming gas, as discussed in Section 4.3.
Finally, the right panel shows the relation of the SFR with the dense, cold gas. As expected, we find a good correlation especially at higher densities. We compare with the linear relation between SFR and H2 found by Bigiel et al. (2008) and extrapolated to lower column densities and find good agreement. That the agreement extends to the normalisation of the relation may be fortuitous. The linearity of the relation is however not a coincidence, as many other studies confirm the existence of a linear relation between dense gas and star formation (Lada et al., 2010; Elia et al., 2022).
@@ 6.3 Current models for the break in the KS-relation
The model of Ostriker et al. (2010) emphasizes the role played by local thermal and vertical mechanical equilibrium in the galactic disk. In particular, a transition in the KS relation is predicted around $\simeq 10$ $\mathrm{M}_{\odot}$ pc-2 as a consequence of the gravity being primarily dominated by the stellar potential or the self-gravity of the gas (see for instance their Figure 2).
The other idea is that the location of the knee is associated to the phase change from HI to H2, instead of from WMN to CNM as proposed in this work (Schaye, 2004; Krumholz et al., 2009b; Krumholz, 2013). This idea is supported by the data from observations which reveals that the SFR is linearly proportional to H2. It also relies on the assumption that star formation occurs in molecular gas exclusively.
@@
Another important aspect that we are not addressing here is the role of the metallicity. It is now well established that metallicity has a significant role on the HI component in galaxies (Wong et al., 2013; Schruba et al., 2018) and does affect the molecular gas content. The possible role metallicity has on the SFR is a bit more controversial. Some studies (e.g. Jameson et al., 2016) found possible significant dependences of the SFR with metallicities in agreement with the predictions of the models presented by Ostriker et al. (2010) and Krumholz (2013) Other studies (Cormier et al., 2014; Bothwell et al., 2016) found that whereas metallicities has a strong influence on the HI and H2 contents, it has a weak impact on the star formation rate itself. This latter conclusion is also met by the numerical simulations of molecular clouds performed by Glover & Clark (2012) where it has been found that the star formation rate would vary by a factor of only 2 when the metallicity was changed by a factor of a hundred.
@@ 7 Model for turbulence induced CNM formation
While many numerical studies have investigated structure formation in a turbulent multi-phase ISM (Audit & Hennebelle, 2005; Hennebelle et al., 2007; Walch et al., 2011; Seifried et al., 2011; Saury et al., 2014; Walch et al., 2015; Kim & Ostriker, 2017; Colman et al., 2024), an analytical model that described the interaction of turbulence with the thermal state of the gas is still lacking. Here, we propose such a model with the aim of predicting the CNM fraction. The general idea is to integrate the mass-weighted density PDF of a turbulent two-phase ISM above a threshold density derived from the temperature-density curve predicted by the cooling/heating model.
@@ 7.1 General assumptions
We do not take into account that CNM formation will lead to star formation which is associated with an increase in UV field strength, ionisation and heating (Ostriker et al., 2010). This could potentially contribute to self-regulating the CNM fraction.

### v3

@@ 2.6 Simulation suite with constant UV field
The parameters of the first set of simulations are summarised in Tables 1 and the final snapshot of each simulation is displayed in Figure 3. We can see a wide variety in terms of structure. The low-$\Sigma$ runs (top panels), especially the ones with not too weak driving, have a prominent diffuse gas component with a thin disk of denser material in which stars form. The high-$\Sigma$ runs (bottom panels) are more filamentary. Increasing the driving strength leads to more large-scale structure, seen as large overdensities and voids (colman22).
@@ 3.1 Star formation history
The qualitative findings described here for modest column densities are different from the results obtained for columns densities above 20 $\mathrm{M}_{\odot}$$\mathrm{pc}^{-2}$, where large-scale driving was found to always reduces the SFR (Brucy_et_al2020; Brucy_et_al2023). This indicates that the relative importance of the processes involved in star formation varies with the column density of the region.
@@ 4.1 Gas phases
In what follows, we will adopt the following definitions for the gas phases. Neutral gas with a temperature below 250 K is considered to be CNM. This threshold temperature is in agreement with the one used in observations (McClure-Griffiths_et_al2023). Note that this definition also includes molecular hydrogen, which is expected to form from cold gas at high densities but whose formation is not tracked explicitly in these simulations. For the lower bound on the temperature of the WNM we choose 3000 K, a value slightly below what is proposed in McClure-Griffiths_et_al2023, to include our lower temperature WNM branch. The LNM is then everything with temperatures in between these thresholds. Our boundaries are shown as dotted lines in Figure 5.
@@ 4.3 Dense gas
We further quantify the efficiency of the formation of dense gas in cold clouds in Figure 8, where we plot the ratio between the amount of CNM versus the amount of gas with densities greater than 100 cm-3. For a point in the bottom left part of the diagram, the amount of cold material is small and the conversion to high density gas inefficient. If a point is in he top part of the diagram, it means the conversion to dense gas is efficient. In the low-$\Sigma$ regime the conversion from cold to dense gas occurs at a roughly equal, albeit low, efficiency which is independent of the level of externally driven turbulence. While the amount of cold HI is similar for all high-$\Sigma$ runs, there is about a factor 2 to 3 less dense gas for the strong driving case. This brings the efficiency to the levels of the lower column density simulations. Strong driving hinders the creation of dense self-gravitating clumps. This is in line with predictions by analytical theories of gravo-turbulent fragmentation (Hennebelle_et_al2024).
@@ 4.4 A dual role of turbulence on the thermal state?
Here we demonstrated that turbulence possibly has a second, thermal, dual nature which typically occurs at larger scales. Hennebelle_Perault1999 already showed that turbulence can create cold condensations in the ISM. Here, we saw that it acts as a perturbing force, promoting phase change and moving the gas away from thermal equilibrium, seen by a consistent increase in LNM fraction (Figure 6). The duality then lays in the creation versus destruction of CNM. Our simulation results seem to indicate turbulence has a stronger impact on the WNM-to-CNM conversion than on the reverse conversion. In fact, there is no indication that turbulence is able to bring the gas back into the WNM state, as the WNM fraction decreases and more gas is in the unstable regime. However, we have to keep in mind that turbulence is not only driven by the external large-scale force that is applied during the simulation. Stellar feedback also drives turbulence. For the set of high-$\Sigma$ simulations, the turbulent energy injection by SN is either stronger or comparable to that of the large-scale driving, except for the run with strong driving (Figure 1). Furthermore, the SFR in those three runs is similar, so only in the case of strong driving would we be able to directly link a change in the gas-phase fractions to an excess of turbulence. If the level of turbulence is a determining factor for CNM formation, we would expect the three other simulations to have similar CNM fractions, as is indeed the case. Possibly the stellar feedback self-regulated the WNM and CNM fractions. Controlled simulation of a multi-phase ISM without stellar feedback, could help quantify the dual role of turbulence on the CNM fraction further. Note that the role of feedback is more complex than just injecting turbulence, as massive stars emit UV radiation which contributes directly to the heating of the ISM.
@@ 4.5 Caveats of a uniform UV background
The transition point between WNM and CNM depends on the UV radiation field, metallicity and hydrogen ionisation rate (Wolfire_et_al2003). While our simulations do track the local UV photon field which is used to determine the ionisation of the gas, this information is not used in the cooling model where a uniform background UV field fixed to the solar neighbourhood value $G_{0}$ is assumed. Observations in the Milky-Way and nearby galaxies show that the UV field strength decreases as a function of galactic radius. The UV field in our low-$\Sigma$ runs may thus be too high compared to real low density regions in galaxies. A lower UV field implies less heating which would lead to naturally higher CNM fraction. Furthermore, in the majority of environments we expect the UV field to be directly coupled to the formation of new stars, especially massive ones. In the next section, we present the result from simulations where the uniform UV background is proportional to the SFR according to the recipe of Ostriker_et_al2010. We will see that this indeed result in more structure formation and enhanced SFRs compared to the low-$\Sigma$ runs presented in this section. It remains however unclear how local variations in the UV field may affect the overall CNM fraction. While we demonstrated a clear influence of turbulence under a fixed UV field, in practice regions with conditions similar to those in the low-$\Sigma$ runs may be rare.
@@ 5 Variable UV background
5.1 Simulation suite with variable UV background
@@
A key difference is the UV background, which is variable in time and proportional to the SFR measured in the computational domain. Indeed, we expect the UV field to be proportional to the SFR (Ostriker_et_al2010). The mean FUV density relative to the solar neighbourhood value $G^{\prime}_{0}$ is
@@
Additionally, we set a minimum UV field of 0.1 $G_{0}$. We use the implementation of Brucy_et_al2023.
@@ 5.4 Velocity dispersion
Figure 11 displays the velocity dispersion as a function of time in the eight simulations of Table 2. A comparison with the runs presented in Figure 2 is particularly enlightening. Except for the two cases with the strongest driving, the velocity dispersions in this set of simulations reach about 11-13 km s-1, irrespective of the forcing strength and initial column density. In contrast, when the UV background is kept constant, the runs with 10.2 M pc-2 present significantly lower velocity dispersion ranging from 2 km s-1 when no turbulent driving is applied, to 7 km s-1 when turbulent forcing operates. This stems from the much higher star formation rate that is achieved in the simulation with variable UV background. The stellar feedback, in particular the number of supernova explosions, is one to two orders of magnitude larger than in the case of constant UV background.
@@ 5.5 Constant or variable UV background?
Although a UV background which is local and proportional to the SFR may be assumed to be the norm, there are situations where a constant UV field may be a good approximation. Let us recall that the mean free path of the UV photons in an ISM with a mean density of 0.3 cm-3 is about 1 kpc. This becomes even larger when the metallicity is subsolar. This implies that for instance in the outskirts of galaxies, the ISM is heated by UV photons which have been emitted at several kpc distance. Accumulated over the galaxy and in a time-span considered by our simulations, the background can then be considered fairly constant.
@@
Figure 12: Relations of the SFR with column density for different gas reservoirs. Top: simulation set with constant UV (Table 1), bottom: set with variable UV (Table 2). The SFR is averaged over a time scale of 10 Myr after (full circles) or before (empty circles) the time of the snapshot at which the gas column densities have been measured. In the left panels, the contours sketch the location of the data point from Bigiel_et_al2008: the outer contour tracers roughly all points, while inner one corresponds roughly to their orange contours denoting 2 independent data point per 0.05 dex. In the middle panels, we show a power-law with index 1.4 to guide the eye. In the right panel, we show the linear relation found by Bigiel_et_al2008 between the SFR and the column density of H2 gas, and extrapolate to lower densities. The simulation data is colour-coded as in Figures 6 and 10.
@@ 6.1 SFR versus different gas phases
In the left panels of Figure 12, we show the total neutral gas versus the SFR. We compare with the observational results from Bigiel_et_al2008, who studied the KS-relation in a sample of 18 galaxies. Their results for HI + H2 are shown as contours, where the outer contour sketches the full extend of their data points and the inner contour indicate where most of the measurements are situated. The observations are characterised by a break at 9 $\mathrm{M}_{\odot}$pc-2. For higher column densities, a power-law relation with index 1.4 is found, while for lower column densities the relation is steeper with a lot of scatter (Kennicutt&Evans2012). The results from the high-$\Sigma$ runs with constant UV field (top panels) fall in the observed range, though the SFR might be a bit high compared to the majority of the observations. Possibly this is because we do not include early feedback in the form of stellar winds and jets, which are expected to reduce the SFR by a factor 2 (Rathjen_et_al2021; Verliat_et_al2022). At low column density, we reproduce the large scatter. The simulations without driving produce SFRs which are below the observed range. This could mean that the environment modelled by this simulation is unlikely to occur in real galaxies. Indeed, large-scale dynamic effects such as rotation or infall would provide a minimal level of turbulence driving. The simulations with strongest driving (orange and green points) generally agree best with the observations.
@@
The middle panels shows the relation with cold gas, that is CNM and cold dense gas which would be molecular. A positive correlation now appears at low column densities. These findings are in line with the results from Smith_at_al2023 who studied the distribution of CNM in their full galaxy simulation. Overall, the relation between the column density of cold gas and the SFR exhibits a scatter of about one order of magnitude around a power-law with index 1.4. This is the same index as found for the relation with total gas, above a column density of 20 $\mathrm{M}_{\odot}$ pc-2. This leads us to the conclusion that the scatter in the observations of the KS-relation at low column densities, as shown in the left panel, is likely due to regional differences in the CNM fraction. Following this line of thought, it implies that at higher column densities the CNM fraction on a scale of 1 kpc does not vary significantly between regions. Indeed, in the simulations we find it to be around 40 % independent of the external driving strength.
Also for the simulations with variable UV background in the bottom panel, we find a similar correlation between the SFR and the cold gas. Again, the strong driving run deviates in behaviour: there is a large variation in SFR for similar column densities. The large velocity dispersions in these runs (see Figure 11), may prevent the formation of sink-forming gas, as discussed in Section 4.3.
Finally, the right panel shows the relation of the SFR with the dense, cold gas. As expected, we find a good correlation especially at higher densities. We compare with the linear relation between SFR and H2 found by Bigiel_et_al2008 and extrapolated to lower column densities and find good agreement. That the agreement extends to the normalisation of the relation may be fortuitous. The linearity of the relation is however not a coincidence, as many other studies confirm the existence of a linear relation between dense gas and star formation (Lada_et_al2010; Elia_et_al2022).
@@ 6.3 Current models for the break in the KS-relation
The model of Ostriker_et_al2010 emphasizes the role played by local thermal and vertical mechanical equilibrium in the galactic disk. In particular, a transition in the KS relation is predicted around $\simeq 10$ $\mathrm{M}_{\odot}$ pc-2 as a consequence of the gravity being primarily dominated by the stellar potential or the self-gravity of the gas (see for instance their Figure 2).
The other idea is that the location of the knee is associated to the phase change from HI to H2, instead of from WMN to CNM as proposed in this work (Schaye2004; Krumholz_et_al2009b; krumholz2013). This idea is supported by the data from observations which reveals that the SFR is linearly proportional to H2. It also relies on the assumption that star formation occurs in molecular gas exclusively.
@@
Another important aspect that we are not addressing here is the role of the metallicity. It is now well established that metallicity has a significant role on the HI component in galaxies (wong2013; schruba2018) and does affect the molecular gas content. The possible role metallicity has on the SFR is a bit more controversial. Some studies (e.g. jameson2016) found possible significant dependences of the SFR with metallicities in agreement with the predictions of the models presented by Ostriker_et_al2010 and krumholz2013 Other studies (cormier2014; bothwell2016) found that whereas metallicities has a strong influence on the HI and H2 contents, it has a weak impact on the star formation rate itself. This latter conclusion is also met by the numerical simulations of molecular clouds performed by Glover2012 where it has been found that the star formation rate would vary by a factor of only 2 when the metallicity was changed by a factor of a hundred.
@@ 7 Model for turbulence induced CNM formation
While many numerical studies have investigated structure formation in a turbulent multi-phase ISM (Audit_Hennebelle2005; Hennebelle_et_al2007; Walch_et_al2011; Seifried_et_al2011; Saury_et_al2014; Walch_et_al2015; Kim_Ostriker2017; colman24), an analytical model that described the interaction of turbulence with the thermal state of the gas is still lacking. Here, we propose such a model with the aim of predicting the CNM fraction. The general idea is to integrate the mass-weighted density PDF of a turbulent two-phase ISM above a threshold density derived from the temperature-density curve predicted by the cooling/heating model.
@@ 7.1 General assumptions
We do not take into account that CNM formation will lead to star formation which is associated with an increase in UV field strength, ionisation and heating (Ostriker_et_al2010). This could potentially contribute to self-regulating the CNM fraction.